# Enhancing Variable Selection in Large-scale Logistic Regression: Leveraging Manual Labeling with Beneficial Noise

### v1

@@ Abstract
In large-scale supervised learning, penalized logistic regression (PLR) effectively addresses the overfitting problem by introducing regularization terms yet its performance still depends on efficient variable selection strategies. This paper theoretically demonstrates that label noise stemming from manual labeling, which is solely related to classification difficulty, represents a type of beneficial noise for variable selection in PLR. This benefit is reflected in a more accurate estimation of the selected non-zero coefficients when compared with the case where only truth labels are used. Under large-scale settings, the sample size for PLR can become very large, making it infeasible to store on a single machine. In such cases, distributed computing methods are required to handle PLR model with manual labeling. This paper presents a partition-insensitive parallel algorithm founded on the ADMM (alternating direction method of multipliers) algorithm to address PLR by incorporating manual labeling. The partition insensitivity of the proposed algorithm refers to the fact that the solutions obtained by the algorithm will not change with the distributed storage of data. In addition, the algorithm has global convergence and a sublinear convergence rate. Experimental results indicate that, as compared with traditional variable selection classification techniques, the PLR with manually-labeled noisy data achieves higher estimation and classification accuracy across multiple large-scale datasets.
Index Terms:
@@
Beneficial noise; Massive data; Logistic regression; Variable selection; Distributed computing.
I Introduction
@@
Label noise is prevalent in classification problems within statistics [1, 2, 3] and machine learning [4, 5, 6, 7]. A common approach focuses on developing methods to maintain the efficiency of classifiers in the presence of label noise. Frenay and Verleysen [8] provided overviews of some works on this problem, and categorized the methods into four types: label noise-robust methods, data cleaning methods, probabilistic label noise-tolerant methods and model-based label noise-tolerant methods. However, all these methods treat noise as either harmful or a lack of statistical information.
@@
Recently, Li [9] pointed out that not all noise is harmful; certain types of noise can exhibit beneficial characteristics in specific tasks. Almost simultaneously, similar conclusions have emerged in the field of statistical learning regarding classification problems, see [10, 11, 12]. Inspired by these works, this paper investigates whether, in the context of large-scale (big) data, there exists a type of beneficial noise that can improve the variable selection performance of PLR in [13]. Variable selection refers to the process of choosing a subset of the most important variables from a large set of candidate variables when building a statistical model, to improve predictive performance and interpretability. By excluding irrelevant or redundant variables, variable selection helps reduce the complexity of a logistic regression model, prevent overfitting, and enhance stability and interpretability.
The PLR loss function is given by:
$L(\bm{\beta})=$ $-\sum_{i=1}^{n}\left[y_{i}\log({p}(\bm{x}_{i},\bm{\beta}))\right.$ (1)
$\left.+(1-y_{i})\log(1-{p}(\bm{x}_{i},\bm{\beta}))\right]$
$+\sum_{j=1}^{d}P_{\lambda}(|\beta_{j}|),$
@@
${p}(\bm{x}_{i},\bm{\beta})=\frac{1}{1+\exp({-\bm{\beta}^{\top}\bm{x}_{i})}}$ (2)
is the predicted probability for the $i$-th observation, $y_{i}\in\{0,1\}$ is the observed label, $\bm{x}_{i}$ is the feature vector, $\bm{\beta}$ is the vector of regression coefficients, and $P_{\lambda}(|\beta_{j}|)$ is the regularization term with parameter $\lambda$. A popular regularization term with variable selection ability is LASSO (least absolute shrinkage and selection operator), also known as $\ell_{1}$ regularization ($\lambda\sum_{j=1}^{d}|\beta_{j}|$). However, many studies [13, 14, 15] have shown that under the assumption of sparse coefficients (where most coefficients are zero and only a few are non-zero), the LASSO regularization term does not possess variable selection consistency. In other words, the estimator of LASSO-PLR in (1) cannot guarantee that the true non-zero coefficients will be estimated as non-zero. To ensure the consistency of variable selection in PLR, an effective alternative is the adaptive (weighted) LASSO (ALASSO [13]) or some nonconvex regularization terms in [15]. In this paper, ALASSO serves as the primary regularization term.
In supervised learning, manual labeling of training samples is a common practice [8], especially when dealing with large datasets where some labels are missing or are not easily obtained. When the labeling task is inherently challenging, the provided labels may not always align with the true labels, thus introducing label noise into the training dataset (that is, label noise comes from classification difficulty). Moreover, when multiple experts manually label the same instance, it is possible for them to assign different category labels, a scenario of label noise that is also considered in this study. To facilitate the use of statistical analysis tools in examining the impact of this noise, we adopt the strategy from [11], where manual labels are randomly sampled based on the posterior probabilities of class membership. In this setting, the noise generated by the label is only related to the difficulty of classification. A more relaxed version involves using a random effects model (a.k.a. finite-mixture model in [16, 17]) that generates manual labels based on imperfect approximations of the posterior probabilities of class membership.
@@
    The first contribution of this paper is to demonstrate that in large-scale PLR, the noise in manual labels (which is solely related to the difficulty of classification) can offer statistical information that is precisely identical to that of the truth label. In other word, it is beneficial to the variable selection performance of PLR.
@@
    The second contribution of this paper is to develop a partition-insensitive parallel ADMM algorithm for solving large-scale PLR with manual labels data stored in a distributed manner. The most prominent feature of this distributed algorithm is its insensitivity to data partitioning. In distributed computing, it means that no matter how the data is stored, regardless of the number of machines the data is stored on and the amount of data each machine holds, the iterative solutions obtained by the algorithm stay consistent. This feature is of great significance for improving result predictability, system scalability, and computational efficiency. Moreover, this solution is exactly the same as the one obtained when a single machine processes the entire dataset. In theory, the distributed algorithm can be proven to have a sublinear convergence rate.
The paper is structured as follows. Section II offers preliminaries and a literature review. Section III presents the variable selection properties and asymptotic relative efficiency of PLR with beneficial noise in manual labels. Section IV proposes a new parallel algorithm for PLR under manual labels. Section V conducts numerical experiments on real datasets. Section VI concludes the paper and suggests future research directions. Supplementary experiments and technical proofs are in the online appendix.
II Preliminaries and Literature Reviews
II-A Penalized Logistic Regression
@@
$f(y\mid\bm{x},\bm{\beta})=\left({p}(\bm{x},\bm{\beta})\right)^{y}\left(1-{p}(\bm{x},\bm{\beta})\right)^{1-y},$ (3)
where ${p}(\cdot)$ is defined in (2). The Fisher information matrix is defined as
$I(\bm{\beta})=\mathbb{E}_{\bm{x}}\left[\bm{x}\bm{x}^{\top}{p}(\bm{x},\bm{\beta})(1-{p}(\bm{x},\bm{\beta}))\right].$ (4)
The dataset $\{\bm{x}_{i},y_{i}\}_{i=1}^{n}$ consists of independently-collected observation samples regarding $\{\bm{x},y\}$. Taking the negative log empirical likelihood function of the collected dataset yields the logistic loss function (cross-entropy loss function), which is the first term in (1).
II-A1 Sparsity and Variable Selection
@@
The method of selecting non-zero coefficients is called variable selection. In logistic regression, a commonly used variable selection method is to add a regularization term as shown in (1). The estimator of ALASSO-PLR is given as follows:
$\hat{\bm{\beta}}\in\arg\min_{\bm{\beta}}$$\left\{-\sum_{i=1}^{n}\left[y_{i}\bm{\beta}^{\top}\bm{x}_{i}-\log(1+e^{\bm{\beta}^{\top}\bm{x}_{i}})\right]\right.$
$\left.\ \ +\sum_{j=1}^{d}{\lambda_{j}}|\beta_{j}|\right\}.$ (5)
The first term is derived from substituting (2) into the first term of (1). The second term is the ALASSO regularization term first proposed by [13]. Specifically, the single regularization parameter $\lambda$ in LASSO is replaced by parameter vector $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{d})^{\top}$. Here, $\lambda_{j}$ varies with the absolute value of $\beta_{j}$. Generally speaking, the larger $|\beta_{j}|$ is, the smaller $\lambda_{j}$ will be. Zou [13] put forward a weighted strategy, which is expressed as
$\bm{\lambda}=\frac{\lambda}{|\bm{\beta}^{{}^{\prime}}|^{\gamma}}.$ (6)
@@
$\|{\bm{\beta}}^{{}^{\prime}}-\bm{\beta}^{*}\|_{2}=O_{P}(n^{-1/2}).$ (7)
A consistent estimator generally produces precise coefficient estimates, especially in the big-data scenario (i.e., when $n$ is extremely large). Subsequently, the weighted strategy presented in (6) can assign small regularization parameters to coefficients with large absolute values and large regularization parameters to those with small absolute values. A natural choice for $\bm{\beta}^{{}^{\prime}}$ is the ordinary logistic regression estimator. Unless otherwise specified below, $\bm{\beta}^{{}^{\prime}}$ will denote the ordinary logistic regression estimator.
@@ II-A2 Oracle Property
[…]
@@
The importance of the oracle property has been emphasized in the literature [18, 19], with the argument that an effective procedure should ideally possess these characteristics.
@@
These two properties are not possessed by ordinary logistic regression and LASSO-PLR in general situations. Specifically, the ordinary logistic regression estimator $\bm{\beta}^{{}^{\prime}}$ has $\sqrt{n}$-consistency in (7) and the following asymptotic distribution
[…]
@@
However, its coefficients do not have sparsity, that is, the estimated coefficients are not exactly zero. In other words, although the estimated values of the coefficients that are truly zero are small, they are not absolutely zero. Therefore, $\bm{\beta}^{{}^{\prime}}$ without sparsity can not have the above two properties. On the other hand, the LASSO-PLR estimator can achieve sparsity. However, since the LASSO regularization applies the same regularization parameter to all estimation coefficients, it will inevitably result in an inability to accurately select non-zero coefficients or cause a significant bias. In detail, when $\lambda$ is not large enough, the LASSO-PLR estimator may fail to shrink certain irrelevant variables to zero, which means it does not satisfy the first property. As the first property is a prerequisite for the second one, the LASSO-PLR estimator (with a not large enough $\lambda$) does not satisfy the second property either. When $\lambda$ is large, the LASSO-PLR estimator will shrink all irrelevant coefficients to zero. Nevertheless, it will also shrink some coefficients with relatively small absolute values to zero while simultaneously over-compressing other coefficients, leading to significant and non-negligible estimation errors. As a result, it does not possess the two properties. The idea of applying adaptive regularization terms in (6) to coefficients in ALASSO-PLR is to address this drawback of LASSO-PLR.
II-B Manual Labeling
@@
The issue of noise in manual labels in supervised learning is an important research area. Early methods treated the truth labels as latent variables and estimated them based on the observed noisy labels [20]. In recent years, research has shifted its focus to constructing generative models for noisy labels to train classifiers [21, 22, 23, 24]. In the analysis of crowdsourced labels, methods have been developed to filter malicious annotators and rank annotators according to their domain expertise [25, 26]. It is commonly assumed that annotators have comparable skills and no malicious behavior, which is reasonable in medical research but may not hold in low-threshold crowdsourcing applications.
@@
Research often assumes that label noise is uniformly distributed within classes [21, 24], but some studies have considered feature-dependent noise [23]. Under fixed class-conditional noise, logistic regression may be inconsistent [27], while Song et al. [28] showed that under class-conditional uniform noise, noisy labels can be regarded as responses in a modified generalized linear model. Moreover, the phenomenon that label noise is more concentrated around classification decision boundaries has also been explored in robust estimation [29] , but there is limited research on relevant statistical modeling and quantification of information loss. In practice, information loss, human errors, and data entry mistakes may lead to label noise. To simplify theoretical analysis, the authors in [11] commonly assumed that label errors stem solely from classification difficulty, and a finite-mixture model was then employed to relax this assumption. In this paper, we also follow this approach.
@@
In [11], the authors indicated that when the source of label noise is only related to classification difficulty, this noise does not cause loss of statistical information and is therefore beneficial. Here, we apply their theoretical results to the PLR scenario. When discussing the variable selection properties in Section II-A, it is necessary to assume that $\{(\bm{x}_{i},{y}_{i})\}_{i=1}^{n}$ are independent and identically distributed. Therefore, we only need to discuss a single sample $(\bm{x}_{i},{y}_{i})$.
II-B1 Beneficial Label Noise

### v2

# Leveraging Noisy Manual Labels as Useful Information: An Information Fusion Approach for Enhanced Variable Selection in Penalized Logistic Regression

@@ Abstract
In large-scale supervised learning, penalized logistic regression (PLR) effectively mitigates overfitting through regularization, yet its performance critically depends on robust variable selection. This paper demonstrates that label noise introduced during manual annotation, often dismissed as a mere artifact, can serve as a valuable source of information to enhance variable selection in PLR. We theoretically show that such noise, intrinsically linked to classification difficulty, helps refine the estimation of non-zero coefficients compared to using only ground truth labels, effectively turning a common imperfection into a useful information resource. To efficiently leverage this form of information fusion in large-scale settings where data cannot be stored on a single machine, we propose a novel partition insensitive parallel algorithm based on the alternating direction method of multipliers (ADMM). Our method ensures that the solution remains invariant to how data is distributed across workers, a key property for reproducible and stable distributed learning, while guaranteeing global convergence at a sublinear rate. Extensive experiments on multiple large-scale datasets show that the proposed approach consistently outperforms conventional variable selection techniques in both estimation accuracy and classification performance, affirming the value of intentionally fusing noisy manual labels into the learning process.
keywords:
@@
Distributed computing , Information fusion , Logistic regression , Massive data , Noisy manual labels , Variable selection
[…]
organization=School of Mathematics and Statistics, Yunnan University,addressline=No.55 Daxuecheng South Rd., Shapingba District, city=Chongqing, postcode=401331, state=Chongqing, country=China
[…]
organization=Department of Statistics and Data Science, Southern University of Science and Technology,addressline=No. 1088 Xueyuan Avenue, Nanshan District, city=Shenzhen, postcode=518055, state=Guangdong, country=China
1 Introduction
@@
Label noise is prevalent in classification problems within statistics [1] and machine learning [2]. A common approach focuses on developing methods to maintain the efficiency of classifiers in the presence of label noise. Frenay and Verleysen [3] provided overviews of some works on this problem, and categorized the methods into four types: label noise-robust methods, data cleaning methods, probabilistic label noise-tolerant methods and model-based label noise-tolerant methods. However, all these methods treat noise as either harmful or a lack of statistical information.
@@
Recently, Li [4] pointed out that not all noise is harmful; certain types of noise can exhibit useful characteristics in specific tasks. Almost simultaneously, similar findings on classification problems emerged in statistical learning. For example, [5] covered k-nearest neighbour, support vector machine and linear discriminant analysis, [6] dealt with logistic regression, and [7] focused on the logistic regression and support vector machine. Inspired by these works, this paper delves into the potential of leveraging noisy information fusion in the context of large-scale (big) data. Specifically, it investigates whether there exists a type of useful noise within this information-fusion scenario that can improve the variable selection performance of PLR in [8]. Variable selection refers to the process of choosing a subset of the most important variables from a large set of candidate variables when building a statistical model, to improve predictive performance and interpretability. By excluding irrelevant or redundant variables, variable selection helps reduce the complexity of a logistic regression model, prevent overfitting, and enhance stability and interpretability.
The PLR loss function in [9] is given by:
$L(\bm{\beta})=-\sum_{i=1}^{n}\left[y_{i}\log({p}(\bm{x}_{i},\bm{\beta}))+(1-y_{i})\log(1-{p}(\bm{x}_{i},\bm{\beta}))\right]+\sum_{j=1}^{d}P_{\lambda}(|\beta_{j}|),$ (1.1)
@@
${p}(\bm{x}_{i},\bm{\beta})=\frac{1}{1+\exp({-\bm{\beta}^{\top}\bm{x}_{i})}}$ (1.2)
is the predicted probability for the $i$-th observation, $y_{i}\in\{0,1\}$ is the observed label, $\bm{x}_{i}$ is the feature vector, $\bm{\beta}$ is the vector of regression coefficients, and $P_{\lambda}(|\beta_{j}|)$ is the regularization term with parameter $\lambda$. A popular regularization term with variable selection ability is LASSO (least absolute shrinkage and selection operator), also known as $\ell_{1}$ regularization ($\lambda\sum_{j=1}^{d}|\beta_{j}|$). However, many studies (see [8] and [10]) have shown that under the assumption of sparse coefficients (where most coefficients are zero and only a few are non-zero), the LASSO regularization term does not possess variable selection consistency. In other words, the estimator of LASSO-PLR in (1.1) cannot guarantee that the true non-zero coefficients will be estimated as non-zero. To ensure the consistency of variable selection in PLR, an effective alternative is the adaptive (weighted) LASSO (ALASSO [8]) or some nonconvex regularization terms in [10]. Due to space limitations in this paper, ALASSO serves as the primary regularization term.
In supervised learning, manual labeling of training samples is a common practice [3], especially when dealing with large datasets where some labels are missing or are not easily obtained. When the labeling task is inherently challenging, the provided labels may not always align with the true labels, thus introducing label noise into the training dataset (that is, label noise comes from classification difficulty). Moreover, when multiple experts manually label the same instance, it is possible for them to assign different category labels, a scenario of label noise that is also considered in this study. To facilitate the use of statistical analysis tools in examining the impact of this noise, we adopt the strategy from [6], where manual labels are randomly sampled based on the posterior probabilities of class membership. In this setting, the noise generated by the label is only related to the difficulty of classification. A more relaxed version involves using a random effects model (a.k.a. finite-mixture model in [11] and [12]) that generates manual labels based on imperfect approximations of the posterior probabilities of class membership.
@@
    The first contribution of this paper is to demonstrate that in large-scale PLR, the noise in manual labels (which is solely related to the difficulty of classification) can offer statistical information that is precisely identical to that of the truth label. In other word, fusing this noisy information from multiple sources is beneficial for variable selection in PLR.
@@
    The second contribution of this paper is to develop a partition-insensitive parallel ADMM algorithm for solving large-scale PLR with manually-labeled data stored in a distributed manner across multiple machines. The most prominent feature of this distributed algorithm is its insensitivity to data partitioning. In distributed computing, it means that no matter how the data is stored, regardless of the number of machines the data is stored on and the amount of data each machine holds, the iterative solutions obtained by the algorithm stay consistent. This feature is of great significance for improving result predictability, system scalability, and computational efficiency. Moreover, this solution is exactly the same as the one obtained when a single machine processes the entire dataset. In theory, the distributed algorithm can be proven to have a sublinear convergence rate.
The paper is structured as follows. Section 2 offers preliminaries and a literature review. Section 3 presents the variable selection properties and asymptotic relative efficiency of PLR with beneficial noise in manual labels. Section 4 proposes a new parallel algorithm for PLR under manual labels. Section 5 conducts numerical experiments on real datasets. Section 6 concludes the paper and suggests future research directions. Supplementary experiments and technical proofs are in the supplementary material. The R package for implementing the algorithm proposed in this paper and some numerical experiments are available at the following https://github.com/xfwu1016/PLRBN.
2 Preliminaries and Literature Reviews
2.1 Penalized Logistic Regression
@@
$f(y\mid\bm{x},\bm{\beta})=\left({p}(\bm{x},\bm{\beta})\right)^{y}\left(1-{p}(\bm{x},\bm{\beta})\right)^{1-y},$ (2.1)
where ${p}(\cdot)$ is defined in (1.2). The Fisher information matrix is defined as
$I(\bm{\beta})=\mathbb{E}_{\bm{x}}\left[\bm{x}\bm{x}^{\top}{p}(\bm{x},\bm{\beta})(1-{p}(\bm{x},\bm{\beta}))\right].$ (2.2)
The dataset $\{\bm{x}_{i},y_{i}\}_{i=1}^{n}$ consists of independently-collected observation samples regarding $\{\bm{x},y\}$. Taking the negative log empirical likelihood function of the collected dataset yields the logistic loss function (cross-entropy loss function), which is the first term in (1.1).
2.1.1 Sparsity and Variable Selection
@@
The method of selecting non-zero coefficients is called variable selection. In logistic regression, a commonly used variable selection method is to add a regularization term as shown in (1.1). The estimator of ALASSO-PLR is given as follows:
$\hat{\bm{\beta}}\in\arg\min_{\bm{\beta}}\left\{-\sum_{i=1}^{n}\left[y_{i}\bm{\beta}^{\top}\bm{x}_{i}-\log(1+e^{\bm{\beta}^{\top}\bm{x}_{i}})\right]+\sum_{j=1}^{d}{\lambda_{j}}|\beta_{j}|\right\}.$ (2.3)
The first term is derived from substituting (1.2) into the first term of (1.1). The second term is the ALASSO regularization term first proposed by [8]. Specifically, the single regularization parameter $\lambda$ in LASSO is replaced by parameter vector $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{d})^{\top}$. Here, $\lambda_{j}$ varies with the absolute value of $\beta_{j}$. Generally speaking, the larger $|\beta_{j}|$ is, the smaller $\lambda_{j}$ will be. Zou [8] put forward a weighted strategy, which is expressed as
$\bm{\lambda}=\frac{\lambda}{|\bm{\beta}^{{}^{\prime}}|^{\gamma}}.$ (2.4)
@@
$\|{\bm{\beta}}^{{}^{\prime}}-\bm{\beta}^{*}\|_{2}=O_{P}(n^{-1/2}).$ (2.5)
A consistent estimator generally produces precise coefficient estimates, especially in the big-data scenario (i.e., when $n$ is extremely large). Subsequently, the weighted strategy presented in (2.4) can assign small regularization parameters to coefficients with large absolute values and large regularization parameters to those with small absolute values. A natural choice for $\bm{\beta}^{{}^{\prime}}$ is the ordinary logistic regression estimator. Unless otherwise specified below, $\bm{\beta}^{{}^{\prime}}$ will denote the ordinary logistic regression estimator.
@@ 2.1.2 Oracle Property
    Asymptotic normality: It achieves the optimal estimation rate, expressed as $\sqrt{n}(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}_{\mathcal{A}}^{*})\xrightarrow{d}\text{N}(0,I_{\mathcal{A}}^{-1})$, where $\xrightarrow{d}$ denotes convergence in distribution.
@@
The importance of the oracle property has been emphasized in the literature [13] and [14], with the argument that an effective procedure should ideally possess these characteristics.
@@
These two properties are not possessed by ordinary logistic regression and LASSO-PLR in general situations. Specifically, the ordinary logistic regression estimator $\bm{\beta}^{{}^{\prime}}$ has $\sqrt{n}$-consistency in (2.5) and the following asymptotic distribution
$\sqrt{n}({\bm{\beta}}^{{}^{\prime}}-\bm{\beta}^{*})\xrightarrow{d}\text{N}(0,I(\bm{\beta}^{*})^{-1}).$ (2.6)
@@
However, its coefficients do not have sparsity, that is, the estimated coefficients are not exactly zero. In other words, although the estimated values of the coefficients that are truly zero are small, they are not absolutely zero. Therefore, $\bm{\beta}^{{}^{\prime}}$ without sparsity can not have the above two properties. On the other hand, the LASSO-PLR estimator can achieve sparsity. However, since the LASSO regularization applies the same regularization parameter to all estimation coefficients, it will inevitably result in an inability to accurately select non-zero coefficients or cause a significant bias. In detail, when $\lambda$ is not large enough, the LASSO-PLR estimator may fail to shrink certain irrelevant variables to zero, which means it does not satisfy the first property. As the first property is a prerequisite for the second one, the LASSO-PLR estimator (with a not large enough $\lambda$) does not satisfy the second property either. When $\lambda$ is large, the LASSO-PLR estimator will shrink all irrelevant coefficients to zero. Nevertheless, it will also shrink some coefficients with relatively small absolute values to zero while simultaneously over-compressing other coefficients, leading to significant and non-negligible estimation errors. As a result, it does not possess the two properties. The idea of applying adaptive regularization terms in (2.4) to coefficients in ALASSO-PLR is to address this drawback of LASSO-PLR.
2.2 Manual Labeling
@@
The issue of noise in manual labels in supervised learning is an important research area. Early methods treated the truth labels as latent variables and estimated them based on the observed noisy labels [15]. In recent years, research has shifted its focus to constructing generative models for noisy labels to train classifiers, see [16] and [17]. In crowdsourced label analysis, methods like those in [18] and [19] have been developed to filter malicious annotators and rank them by domain expertise. It is commonly assumed that annotators have comparable skills and no malicious behavior, which is reasonable in medical research but may not hold in low-threshold crowdsourcing applications.
@@
Research often assumes that label noise is uniformly distributed within classes (see [20] and [21]), but some studies have considered feature-dependent noise [17]. Under fixed class-conditional noise, logistic regression may be inconsistent [22], while Song et al. [23] showed that under class-conditional uniform noise, noisy labels can be regarded as responses in a modified generalized linear model. Moreover, the phenomenon that label noise is more concentrated around classification decision boundaries has also been explored in robust estimation [24] , but there is limited research on relevant statistical modeling and quantification of information loss. In practice, information loss, human errors, and data entry mistakes may lead to label noise. To simplify theoretical analysis, the authors in [6] commonly assumed that label errors stem solely from classification difficulty, and a finite-mixture model was then employed to relax this assumption. In this paper, we also follow this approach.
@@
In [6], the authors indicated that when the source of label noise is only related to classification difficulty, this noise does not cause loss of statistical information and is therefore beneficial. Here, we apply their theoretical results to the PLR scenario. When discussing the variable selection properties in Section 2.1, it is necessary to assume that $\{(\bm{x}_{i},{y}_{i})\}_{i=1}^{n}$ are independent and identically distributed. Therefore, we only need to discuss a single sample $(\bm{x}_{i},{y}_{i})$.
2.2.1 Useful Label Noise